OpenStudy (hoothoot23):

An equation is shown below: 8x + 2(x – 7) = 7x + 3x – 14 Part A: Solve the equation and write the number of solutions. Show all the steps. (6 points) Part B: Name one property you used to solve this equation.

OpenStudy (hoothoot23):

@alekos

rishavraj (rishavraj):

@hoothoot23 did u try ??? wht do u think would be the first step ??

OpenStudy (alekos):

Didnt someone just solve this for you?

OpenStudy (hoothoot23):

no

rishavraj (rishavraj):

wht do u think would be the first step ??

OpenStudy (hoothoot23):

okay let me fihure out the first step

OpenStudy (alekos):

is this the second of a similar problem you just posted?

OpenStudy (hoothoot23):

8x + 2(x – 7) = 7x + 3x – 14 10x+14=7x+3x-14

OpenStudy (hoothoot23):

and no it isn't

OpenStudy (hoothoot23):

thats what i thought

rishavraj (rishavraj):

10x - 14 = 10x - 14

OpenStudy (hoothoot23):

oh

rishavraj (rishavraj):

@hoothoot23 u hav typed +14 ??

OpenStudy (hoothoot23):

im confused can you explain the steps

rishavraj (rishavraj):

see 8x + 2(x – 7) = 7x + 3x – 14 8x + 2x - 14 = 10x - 14 10x - 14 = 10x - 14 wht do u think the next step sould be?

rishavraj (rishavraj):

@hoothoot23